A sample of nitrogen gas expands in volume from 1.6 L to 5.4 L at constant temperature. Calculate the work done in joules if the gas expands (a) against a vacuum, (b) against a constant pressure of 0.80 atm, and (c) against a constant pressure of 3.7 atm.
1
Expert's answer
2020-09-14T08:22:12-0400

A sample of nitrogen gas expands in volume from 1.6 L to 5.4 L at constant temperature


a) Work done by vaccumn = zero .


b) Work done = p ( v2 - v1 ) = 0.80 ( 5.4 - 1.6 ) atm . liter = 3.04 atm . = ( 3.04 * 101.2 ) j =

307.6 j


c) Work done = p ( v2 - v1 ) .

= 3.7 ( 5.4 - 1.6 ) atm . liter

= 14.06 atm . liter

= 14.06 * 101.2 j .

= 1422.872 j .